# Formatting instructions for NeurIPS 2018

### v1

@@ 1 Introduction
As more machine learning models are adopted in real-world production pipelines, there is an increasing demand for interpretability to obtain key analytical insights which can build trust in AI-driven decision support systems [1, 2, 3]. Such model understanding enables developers to evaluate a model’s reliability, which is crucial for high-stake decision spaces such as politics and healthcare [4, 5, 6].
However, when a machine learning model is updated as new data is incorporated into training, analytical insights often diverge significantly from previous observations [7]. As users develop trust in models over time based on their past performance and model alignment with personal expertise, when insights suddenly shift due to new data update, it can lead to skepticism and hesitation for adoption .
@@
Consider a predictive model used to estimate a patient’s mortality in a hospital: if the model’s insights drastically change with each new iteration, physicians may find it challenging to explain these variations to their patients and may even create legal difficulties if incidents of error arise. Patients rely on doctors’ recommendations, and if the predictions keep fluctuating, it can challenge their confidence in the medical decision-making process [4].
@@
There are multiple avenues to address challenges faced by retraining in the literature, via continual learning [8], online learning [9], as well as many others [10, 11, 12, 13]. However, existing methods focus largely on retaining prior performance on the retrained models in the face of data distributional shifts, or conceptual shifts, and thus do not put an emphasis on the fact that analytical insights: such as the decision path of a prediction, or important features selected by a model, be stable across different retraining iterations. In this work, we demonstrate a direct incorporation of analytical insights into training more stable models and ensure that the output models provide consistent analytical outcomes.
@@
Additionally, retraining should be distinguished from online learning, which aims at continuously updating data and more importantly, incrementally updating the machine learning model instead of retraining the model from scratch with each data batch update (batch learning). One advantage of batch learning is its smaller consumption of computational resources and faster retraining speed since it is done in a periodical instead of a continuous fashion. Furthermore, such a training strategy allows for a more global solution, as retraining involves using the entire database, whereas online learning, often to ensure faster speed, incrementally incorporates only a portion of the original data, and provides more greedy solutions [14].
@@
Recently, another body of literature focusing on model structure stability has been investigated in the operations research and computer science community. Various methods, such as mixed integer optimization, were used in [15, 16] to achieve sequentially stable regression and classification tree models. Other works including [17] also developed methods that produce stable feature interactions in random forest-based models. Our work follows and extends the work of [16] by developing a unifying framework that learns stable structure across multiple different modeling techniques, including regression, classification trees, and gradient-boosted trees.
@@ 1.1 Contributions
We summarize the main contributions of this paper: (i) we develop a general framework aimed at stabilizing machine learning model structures and formulate it as a mixed-integer optimization-based algorithm for retraining machine learning models that improve existing greedy approaches. (ii) we demonstrate a significant boost in model stability with small, computational efficiency, and controlled sacrifice of performance. (iii) we show its applicability in an existing production pipeline used in a large US hospital for mortality prediction.
The rest of the paper is organized as follows. In Section 2, we outline the problem setting and formulation of the retraining setting, in Section 3, we present numerical experimental results on a real-world case study, in 4, we discuss our limitations and conclude our study.
@@ 2 A Methodology for Retraining Machine Learning Models
In this section, we introduce the problem setting of the retraining problem, its formulation as a mixed integer problem, and finally discuss the adaptive training strategies.
@@ 2.1 Problem Setting
Consider an observational dataset composed of triplets $(\boldsymbol{x_{i}},y_{i},t_{i})_{i=1}^{N}$, where $\boldsymbol{x_{i}}\in\mathbb{R}^{p}$ represents the feature vector, $y_{i}\in\mathbb{R}$ is the target variable, and $t_{i}$ is the time at which the data point $i$ was obtained. The dataset is accumulated over time, resulting in a series of batches at times $T_{1}<T_{2}<\ldots<T_{B}$. The dataset available at time $T_{b},\ b\in[B]:=\{1,\ldots,B\},$ is denoted by $D_{b}$ and is defined as:
@@ 2.2 Problem Statement
Minimizing Predictive Error: The first requirement is to ensure that each model $f_{b}$ in the sequence $f_{1},\ldots,f_{B}$ has a low out-of-sample predictive error. Formally, denoting by $\mathcal{D}_{b}$ the distribution of the training data in $D_{b}$, we seek to minimize the expected error for data points $(\boldsymbol{x},y)\sim\mathcal{D}_{b}$:
$\min_{f_{1},\ldots,f_{B}}\sum_{b=1}^{B}\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}_{b}}[\ell(f_{b}(\boldsymbol{x}),y)],$ (1)
where the loss function $\ell(\cdot,\cdot)$ quantifies the prediction error when the predicted value is $f_{b}(\boldsymbol{x})$ and the true value is $y$.
@@ 2.3 A General Formulation
Depending on the type of the model used, solving Problem (4) can be challenging, as we need to consider either an infinite number (i.e. neural network) or an exponentially growing number (with respective to feature size, i.e. decision tree) of possible functions $f_{b}$. This is particularly true for models that are commonly used in high-stakes healthcare applications, e.g., decision trees and variants thereof. Thus, we next develop a restricted formulation that can tractably accommodate any type of machine learning model.
@@ 2.5 Model Distance Calculation
  • For linear/logistic regression models, we compute the distance as $d(f_{i},f_{j})=\|\boldsymbol{\beta}^{i}-\boldsymbol{\beta}^{j}\|_{2}^{2}$, where $\boldsymbol{\beta}^{i}$ and $\boldsymbol{\beta}^{j}$ are the regressions coefficient vectors of $f_{i}$ and $f_{j}$, respectively.
  • For decision tree-based models, we use the decision tree distance metric proposed by [16] that relies on optimally matching the paths in the two trees $\mathbb{T}_{1}$ and $\mathbb{T}_{2}$, denoted as $\mathcal{P}(\mathbb{T}_{1})$ and $\mathcal{P}(\mathbb{T}_{2})$. The distance computation is as follows: first, we calculate the pairwise distance between each path in the first tree and each path in the second tree $d(p,q)$; this pairwise distance takes into account a path weight $w(p)$, defined as the differences in feature ranges in the two paths as well as the class labels associated with them. Then, we determine the overall distance between two decision trees by finding the optimal matching of these paths, a process that can be computed in polynomial time as a variant of the bipartite matching problem. This distance is defined as an integer linear optimization problem, where given decision variables $x_{pq}=\mathbb{I}$(path $p$ in $\mathbb{T}_{1}$ is matched with path $q$ in $\left.\mathbb{T}_{2}\right)$ and $x_{p}=$ $\mathbb{I}$ (path $p$ in $\mathbb{T}_{1}$ is left unmatched), we formulate it as below:
    $d\left(\mathbb{T}_{1},\mathbb{T}_{2}\right)=\min_{\boldsymbol{x}}$ $\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}\sum_{q\in\mathcal{P}\left(\mathbb{T}_{2}\right)}d(p,q)x_{pq}+\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}w(p)x_{p}$
    s.t. $\sum_{q\in\mathcal{P}\left(\mathbb{T}_{2}\right)}x_{pq}+x_{p}=1,\quad\forall p\in\mathcal{P}\left(\mathbb{T}_{1}\right)$
    $\sum_{p\in\mathcal{P}\left(\mathbb{T}_{1}\right)}x_{pq}=1,\quad\forall q\in\mathcal{P}\left(\mathbb{T}_{2}\right)$
    $x_{pq}\in\{0,1\},\quad x_{p}\in\{0,1\},\quad\forall p\in\mathcal{P}\left(\mathbb{T}_{1}\right),\quad\forall q\in\mathcal{P}\left(\mathbb{T}_{2}\right)$
Feature Importance-Based Distances: This approach measures the distance by comparing the models based on their respective feature importances. In particular, denoting by $\boldsymbol{\phi}^{f}\in\mathbb{R}^{p}$ the $p$-dimensional vector of feature importances that model $f$ assigns to each feature, we compute the distance as $d(f_{i},f_{j})=\|\boldsymbol{\phi}^{f_{i}}-\boldsymbol{\phi}^{f_{j}}\|_{2}^{2}$. The advantage of using feature importance-based distances is its versatility to be applied to both interpretable (decision-tree and logistic regression), as well as black-box (gradient boosted tree) models. Concretely, we consider the following two approaches to define feature importance measures:

### v2

# Towards Stable Machine Learning Model Retraining via Slowly Varying Sequences

@@ 1 Introduction
The adoption of machine learning models in real-world production pipelines increases the need for interpretability to obtain analytical insights and build trust in AI-driven decision support systems [1, 2, 3]. Model understanding is crucial for high-stakes decision spaces [4] including healthcare [5, 6] and the social sciences [7]: users develop trust in models over time based on the models’ past performance and alignment with the users’ personal expertise. In practice, machine learning models need to frequently be updated as more data becomes available and is incorporated into training; as a result, the model used in production may diverge over time [8]. Significant shifts in the model’s structure and the resulting analytical insights may may lead to skepticism and hesitation for adoption.
@@
Consider a predictive model used to estimate a patient’s mortality in a hospital: if the model’s insights drastically change with each new iteration, physicians may find it challenging to explain these variations to their patients and may even create legal difficulties if incidents of error arise. Patients rely on doctors’ recommendations and,if the predictions fluctuate, it can challenge their confidence in the medical decision-making process [6].
@@
There are multiple avenues to address challenges faced by retraining in the literature, via continual learning [9], online learning [10], as well as many others [11, 12, 13, 14]. Existing methods focus largely on retaining prior performance on the retrained models in the face of data distributional shifts, or conceptual shifts, and thus do not put an emphasis on the fact that analytical insights: such as the decision path of a prediction, or important features selected by a model, be stable across different retraining iterations. In this work, we demonstrate a direct incorporation of analytical insights into training more stable models and ensure that the output models provide consistent analytical outcomes.
@@
Additionally, retraining should be distinguished from online learning, which aims at continuously updating data and more importantly, incrementally updating the machine learning model instead of retraining the model from scratch with each data batch update (batch learning). One advantage of batch learning is its smaller consumption of computational resources and faster retraining speed since it is done in a periodical instead of in a continuous fashion. Furthermore, such a training strategy allows for a more global solution, as retraining involves using the entire database, whereas online learning, often to ensure faster speed, incrementally incorporates only a portion of the original data, and provides more greedy solutions [15].
@@
Recently, another body of literature focusing on model structure stability has been investigated in the operations research and computer science community. Various methods, such as mixed integer optimization, were used in [16, 17] to achieve sequentially stable regression and classification tree models. Other works including [18] also developed methods that produce stable feature interactions in random forest-based models. Our work follows and extends the work of [17] by developing a unifying, model-agnostic framework that learns stable structure across multiple different modeling techniques, including regression, classification trees, and gradient-boosted trees.
@@ 1.1 Contributions
We summarize the main contributions of this paper: (i) We develop a general framework aimed at stabilizing machine learning model structures upon retraining with new data. We propose an exact mixed-integer optimization formulation and a heuristic restricted formulation for the problem, and develop a practical, model-agnostic algorithm for retraining machine learning models. (ii) We demonstrate, across a variety of models (including logistic regression, classification trees, and boosted trees), a significant boost in model stability with a small and controllable sacrifice of predictive power and computational efficiency compared to existing greedy approaches. (iii) We show our method’s applicability in an existing production pipeline used in a large US hospital for mortality prediction.
The rest of the paper is organized as follows. In Section 2, we outline the problem setting and develop the proposed retraining methodology; in Section 3, we present numerical experimental results on a real-world case study; in 4, we discuss the limitations and conclude our study.
@@ 2 A Methodology for Retraining Machine Learning Models
In this section, we formalize and formulate the machine learning model retraining problem as an optimization problem, we develop a model-agnostic algorithm for retraining machine learning models, and, finally, we discuss adaptive training strategies.
@@ 2.1 Problem Setting
Consider an observational dataset composed of triplets $(\bm{x_{i}},y_{i},t_{i})_{i=1}^{N}$, where $\bm{x_{i}}\in\mathbb{R}^{p}$ represents the feature vector, $y_{i}\in\mathbb{R}$ is the target variable, and $t_{i}$ is the time at which data point $i$ was obtained. The dataset is accumulated over time, resulting in a series of batches at times $T_{1}<T_{2}<\ldots<T_{B}$. The dataset available at time $T_{b},\ b\in[B]:=\{1,\ldots,B\},$ is denoted by $D_{b}$ and is defined as:
@@ 2.2 Problem Statement
Minimizing Predictive Error: The first requirement is to ensure that each model $f_{b}$ in the sequence $f_{1},\ldots,f_{B}$ has a low out-of-sample predictive error. Formally, denoting by $\mathcal{D}_{b}$ the distribution of the training data in $D_{b}$, we seek to minimize the expected error for data points $(\bm{x},y)\sim\mathcal{D}_{b}$:
$\min_{f_{1},\ldots,f_{B}}\sum_{b=1}^{B}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}_{b}}[\ell(f_{b}(\bm{x}),y)],$ (1)
where the loss function $\ell(\cdot,\cdot)$ quantifies the prediction error when the predicted value is $f_{b}(\bm{x})$ and the true value is $y$.
@@ 2.3 A General Formulation
Depending on the type of the model used, solving Problem (4) can be challenging: the number of possible functions $f_{b}$ may grow exponentially with the number of features. This is particularly true for models that are commonly used in high-stakes healthcare applications, including decision trees and variants thereof: considering $p$ binary features and a binary classification task, then there are $2^{2^{p}}$ different ways of labeling all instances, each corresponding to a different underlying boolean function that can be represented as a decision tree. Thus, we next develop a restricted formulation that can tractably accommodate any type of machine learning model.
@@ 2.5 Model Distance Calculation
  • For linear/logistic regression models, we compute the distance as $d(f_{i},f_{j})=\|\bm{\beta}^{i}-\bm{\beta}^{j}\|_{2}^{2}$, where $\bm{\beta}^{i}$ and $\bm{\beta}^{j}$ are the regressions coefficient vectors of $f_{i}$ and $f_{j}$, respectively.
  • For decision tree-based models, we use the decision tree distance metric proposed by [17] that relies on optimally matching the paths in the two trees $\mathbb{T}_{1}$ and $\mathbb{T}_{2}$, denoted as $\mathcal{P}(\mathbb{T}_{1})$ and $\mathcal{P}(\mathbb{T}_{2})$. The distance computation is as follows: first, we calculate the pairwise distance between each path in the first tree and each path in the second tree $d(p,q)$; this pairwise distance takes into account the differences in feature ranges in the two paths as well as the class labels associated with them. Then, we determine the overall distance between two decision trees by finding the optimal matching of these paths, a process that can be computed in polynomial time as a variant of the bipartite matching problem. Formally, assuming without loss of generality that $\mathbbm{T}_{1}$ consists of a larger number of paths, and given decision variables $x_{pq}=\mathbb{I}$(path $p$ in $\mathbb{T}_{1}$ is matched with path $q$ in $\left.\mathbb{T}_{2}\right)$, the distance is defined as the following integer linear optimization problem:
    $\begin{split}d(\mathbbm{T}_{1},\mathbbm{T}_{2})\ =\ \min_{\bm{x}}&\quad\sum_{p\in\mathcal{P}(\mathbb{T}_{1})}\sum_{q\in\mathcal{P}(\mathbb{T}_{2})}d(p,q)x_{pq}\\ \text{s.t.}&\quad\sum_{q\in\mathcal{P}(\mathbb{T}_{2})}x_{pq}=1,\quad\forall p\in\mathcal{P}(\mathbb{T}_{1})\\ &\quad\sum_{p\in\mathcal{P}(\mathbb{T}_{1})}x_{pq}\geq 1,\quad\forall q\in\mathcal{P}(\mathbb{T}_{2})\\ &\quad x_{pq}\in\{0,1\},\quad\forall p\in\mathcal{P}(\mathbb{T}_{1})\ \forall q\in\mathcal{P}(\mathbb{T}_{2}).\end{split}$ (6)
Feature Importance-Based Distances: This approach measures the distance by comparing the models based on their respective feature importances. In particular, denoting by $\bm{\phi}^{f}\in\mathbb{R}^{p}$ the $p$-dimensional vector of feature importances that model $f$ assigns to each feature, we compute the distance as $d(f_{i},f_{j})=\|\bm{\phi}^{f_{i}}-\bm{\phi}^{f_{j}}\|_{2}^{2}$. The advantage of using feature importance-based distances is its versatility to be applied to both interpretable (decision-tree and logistic regression), as well as black-box (gradient boosted tree) models. Concretely, we consider the following two approaches to define feature importance measures: